# ORCID: 0009-0005-8104-2142

### v2

@@ Abstract
This work develops a predictive model to identify potential targets of activist investment funds, which strategically acquire significant corporate stakes to drive operational and strategic improvements and enhance shareholder value. Predicting these targets is crucial for companies to mitigate intervention risks, for activists to select optimal targets, and for investors to capitalize on associated stock price gains. Our analysis utilizes data from the Russell 3000 index from 2016 to 2022. We tested 123 variations of models using different data imputation, oversampling, and machine learning methods, achieving a top AUC-ROC of 0.782. This demonstrates the model’s effectiveness in identifying likely targets of activist funds. We applied the Shapley value method to determine the most influential factors in a company’s susceptibility to activist investment. This interpretative approach provides clear insights into the driving forces behind activist targeting. Our model offers stakeholders a strategic tool for proactive corporate governance and investment strategy, enhancing understanding of the dynamics of activist investing.
@@ 1 Introduction
The participation of activists has been growing recent years, particularly after the financial scandals and the financial crisis of 2007-2008 [12]. In 2018, nearly one campaign launched against a new target every day [19]. Thanks to regulatory reforms that have made shareholder engagement more accessible and weakened corporate defenses, activists can now more easily challenge management to enhance shareholder benefits [9]. The motivations behind activist investing are diverse. Romito identifies four main types of activist investors [27]. Some focus on enhancing returns by advising management to issue debt or redirect free cash flow toward dividends or stock buybacks. Others advocate for merging similar competitors to capitalize on economies of scale. Another group prefers to enhance profitability by separating mismatched divisions within a company. Additionally, some investors aim to streamline operations by maintaining only the most profitable units and spinning off less effective ones. Beyond these financial goals, certain socially responsible funds also promote environmental, social, and governance (ESG) improvements, aiming to foster ethical business practices [4].
@@ 2 Literature Review
Activist funds have long been known to focus on traditional valuation metrics when selecting targets, as revealed by various studies. Pfirrmann and Eichner’s study based on 30 campaigns between 2013 and 2019, reveals a preference for relative valuation over intrinsic valuations, primarily utilizing Enterprise Value/EBITDA (EV/EBITDA) and Price/Earnings (P/E) ratios [24]. Consistent with findings by Brav et al. [8] and Boyson and Mooradian [7], it is observed that firms targeted by activists often exhibit lower valuations compared to their fundamental values. Operationally, these firms show robust cash flows from operations and sales growth rates. Furthermore, their asset returns outperform those of their competitors, despite their stock performance trailing behind the broader market. The analysis also highlights that activist investors typically secure a significant stake in these firms, purchasing between 5
@@ 3 Methodology
Figure 1 provides an overview of the proposed pipeline for activist investment target prediction, consisting of several stages. It begins by collecting data from Russell 3000 companies (2016-2022) and activist campaigns (2011-2023), featuring diverse variables. In preprocessing, non-ratio variables are transformed into ratios and selected variables are converted to percentiles relative to industry peers. Four imputation techniques handle missing data, and four oversampling methods address class imbalance. Six machine learning models are trained using these imputed and balanced datasets. SHAP values are then used to interpret the models, highlighting key variables. Each step of this framework is detailed in the following subsections.
@@ 3.2.2 Data Imputation
Missing data is a common issue in data science, arising from factors like human error or incomplete records, and it can impede analysis. Imputation methods provide a systematic approach to estimate and fill in these gaps, allowing for more robust data analysis without the need to discard incomplete observations. We use four different methods: Mean Imputation, K-nearest neighbours (KNN) imputations (k=5) [5], Multivariate Imputation by Chained Equations using LightGBM (MiceForest) [29], and Generative Adversarial Imputation Networks (GAIN) [31]. For further details of the methods, refer to Appendix B.
@@ 3.4 Model Interpretation Method
In machine learning, there is an inherent trade-off between model complexity and interpretability. Advanced models like neural networks often outperformed simpler alternatives such as linear regression or decision trees in terms of predictive accuracy but typically offered less interpretability. To address the challenge of understanding and interpreting complex models, various methodologies have been developed.

### v3

@@ Abstract
This research presents a predictive model to identify potential targets of activist investment funds—entities that acquire significant corporate stakes to influence strategic and operational decisions, ultimately enhancing shareholder value. Predicting such targets is crucial for companies aiming to mitigate intervention risks, activist funds seeking optimal investments, and investors looking to leverage potential stock price gains. Using data from the Russell 3000 index from 2016 to 2022, we evaluated 123 model configurations incorporating diverse imputation, oversampling, and machine learning techniques. Our best model achieved an AUC-ROC of 0.782, demonstrating its capability to effectively predict activist fund targets. To enhance interpretability, we employed the Shapley value method to identify key factors influencing a company’s likelihood of being targeted, highlighting the dynamic mechanisms underlying activist fund target selection. These insights offer a powerful tool for proactive corporate governance and informed investment strategies, advancing understanding of the mechanisms driving activist investment decisions.
@@ 1 Introduction
The participation of activists has been growing in recent years, particularly after the financial scandals and the financial crisis of 2007–2008 [12]. In 2018, nearly one campaign launched against a new target every day [19]. Thanks to regulatory reforms that have made shareholder engagement more accessible and weakened corporate defenses, activists can now more easily challenge management to enhance shareholder benefits [9]. The motivations behind activist investing are diverse. Romito identifies four main types of activist investors [27]. Some focus on enhancing returns by advising management to issue debt or redirect free cash flow toward dividends or stock buybacks. Others advocate for merging similar competitors to capitalize on economies of scale. Another group prefers to enhance profitability by separating mismatched divisions within a company. Additionally, some investors aim to streamline operations by maintaining only the most profitable units and spinning off less effective ones. Beyond these financial goals, certain socially responsible funds also promote environmental, social, and governance (ESG) improvements, aiming to foster ethical business practices [4].
@@ 2 Literature Review
Activist funds have long been known to focus on traditional valuation metrics when selecting targets, as revealed by various studies. Pfirrmann and Eichner’s study based on 30 campaigns between 2013 and 2019, reveals a preference for relative valuation over intrinsic valuations, primarily utilizing Enterprise Value/EBITDA (EV/EBITDA) and Price/Earnings (P/E) ratios [24]. Consistent with findings by Brav et al. [8] and Boyson and Mooradian [7], it is observed that firms targeted by activists often exhibit lower valuations compared to their fundamental values. Operationally, these firms show robust cash flows from operations and sales growth rates. Furthermore, their asset returns outperform those of their competitors, despite their stock performance trailing behind the broader market. The analysis also highlights that activist investors typically secure a significant stake in these firms, purchasing between 5% to 10% of equity shares on the market once they have identified their targets.
@@ 3 Methodology
Fig. 1 provides an overview of the proposed pipeline for activist investment target prediction, consisting of several stages. It begins by collecting data from Russell 3000 companies (2016-2022) and activist campaigns (2011-2023), featuring diverse variables. In preprocessing, non-ratio variables are transformed into ratios and selected variables are converted to percentiles relative to industry peers. Four imputation techniques handle missing data, and four oversampling methods address class imbalance. Six machine learning models are trained using these imputed and balanced datasets. SHAP values are then used to interpret the models, highlighting key variables. Each step of this framework is detailed in the following subsections.
@@ 3.2.2 Data Imputation
A considerable portion of dataset contains missing values, posing challenges to analysis (see Fig. 4.2.1). Imputation methods estimated and fill in these gaps, enabling robust analysis without discarding incomplete observations. We use four different methods: Mean Imputation, K-nearest neighbors (KNN) imputations (k = 5) [5], Multivariate Imputation by Chained Equations using LightGBM (MiceForest) [29], and Generative Adversarial Imputation Networks (GAIN) [31]. For further details of the methods, refer to Appendix B.
@@ 3.4 Model Interpretation Method
The trade-off between model complexity and interpretability is a key challenge in machine learning. Neural networks and various ensemble models often outperform simpler models like linear regression but fall short in interpretability. This has led to the development of various methodologies.